OpenStudy (anonymous):

The perimeter of a rectangle is 42cm. The length is 3cm greater than the width. Find the width.

OpenStudy (espex):

Recall that P = 2*L + 2*W, and you're given that L = W+3 Rewrite the equation to P = 2(W+3) + 2*W Plug in your values and solve for W. :)